The width of hotographis 4 centimeters more that three tenths of the length. If the width is 13 cm, find the length.
Answer by stormy(4)

I'm assuming you mean "The width of a photograph is..."

Let W = the width and L = the length, both in centimeters.

The width is: " w = "
Four cm more than: " 4 + "
Three-tenths of the length: " 0.3 * L "

Now simply put it all together:

w = 4 + 0.3 * L

Substitute 13 for the width, and solve for L:

13 = 4 + 0.3 * L
13 - 4 = 0.3 * L
9 = 0.3 * L
9 / 0.3 = L
L = 30 cm

To check the solution, go back to the original problem. Three tenths the length is:

0.3 * 30 = 9

Four centimeters more than this amount is:

9 + 4 = 13 cm

This is the width given in the problem, so our answer is correct.